OpenStudy (anonymous):

Find each percent of change. Round answers to the nearest tenth, if necessary. 1. 42 is increased to 63.

OpenStudy (marsxtc):

63-42 = 21 21/42 = .5 *100 = 50%

OpenStudy (anonymous):

I have 150%.

OpenStudy (anonymous):

cant be 150 percent because 100 percent change would be 84